Get Help

Online: Request Help
Phone: (707) 826-HELP (4357)
Email: help@humboldt [dot] edu
In Person: Library 120 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
CMS/Peoplesoft Feb 22 2012 - 18:00 Anyone who uses CFS Production (FCFSPRD) and CFS Data Warehouse application and database services, including process schedulers and integration broker CMS Central will install the following changes (none that affect the functionality for campus end users): 1. BSL_CFS_FS_FS09001 Update to Data Security HD223733 2. BSL_CFS_PO_PO99011 Update to Purchase Order Print FormatHD223987 3. BSL_CFS_PO_ZFS90024 Update to Row-Level Security of Comments Table HD209718 / INC0016910 4. BSL_CFS_AP_ZFS90025 Update to Pay Cycle Pre-Note Creation INC0018757
Other Feb 20 2012 - 09:31 Users of CNRS-file Hardware Failure
Other Feb 19 2012 - 05:00 All Users Due to unexpected issues coming up, the February Maintenance Window has to be extended until 1:00 p.m.
Other Feb 19 2012 - 10:39 All SHC Orchard Harvest Users The server was patched during the February Maintenance Window, and stopped working thereafter. ITS-Sysadmins emailed SHC-EMR-Support@humboldt.edu
OBI Feb 16 2012 - 10:05 Finance data is old or not available investigating.
Campus Web Sites Feb 15 2012 - 13:00 all users accessing www3 server Server out of memory. Rebooted by ITS sysadmins. Still investigating cause.
Wireless Network Feb 14 2012 - 13:25 Wireless users - Science A 5th & 6th Floor UPDATE: ***** WIRELESS IS NOW AVAILABLE ***** A couple of Access Points have gone down, TNS is looking into it. Please check back for updates.
Network Folders Feb 12 2012 - 15:59 All Network Folders Users A CNRS student was running project related processes, that taxed the storage on the main Folders server enough for Samba connections (U: drives, etc.) to timeout. The system has recovered at this point. A KBox ticket for CNRS-ITC has been issued to "talk" to the user.
Campus Data Network Feb 7 2012 - 15:00 Users in Residence Halls using ResnetWireless or a wired port UPDATE: FEB 7TH, 2012 @ 4PM. The unauthorized device has been isolated. --------ORIGINAL MESSAGE-------- A resident in the west canyon buildings is running an unauthorized DHCP server. We are investigating.
Off-Campus Internet Connection Feb 19 2012 - 04:00 N/A Demand for bandwidth continues to grow every day. Our primary edge router processing engine has reached the forwarding rate limit and must be replaced.
Campus Data Network Feb 9 2012 - 16:00 All Students, Faculty, & Staff occupying the University Center #45. As part of the campus ITRP2 network project, network equipment will be replaced every 3-5 years to keep the campus data network modern, secure and robust.
Campus Web Sites Feb 1 2012 - 13:54 All Users trying to connect to Central Web Apache had 170 threads, the server was swapping, and MySQL was not responding. Restarted Apache, and MySQL. ITS Sysadmins are looking into the root cause. It appears, that the buggy /events site was over-taxing MySQL. An known bug-heavy feature on the /events site will be disabled while the entire site is being debugged.
CMS/Peoplesoft Feb 1 2012 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions HUM_AD401, HUM_AD401_SAT_PREFILTER.
Other Jan 31 2012 - 16:45 All VLAB Users The hypervisor serving the Web server for VLAB had to be shut down by ITS Sysadmins due to emergency maintenance. XEN hypervisor OS file system ran out of space.
Other Jan 31 2012 - 16:09 All NOLIJ Users The XEN Hypervisor running NOLIJ has frozen up. ITS-Sysadmins are investigating. ITS Sysadmins had to reboot other XEN cluster nodes to troubleshoot this problem. XEN hypervisor OS file system ran out of space.
Directory Services Jan 31 2012 - 16:09 All Windows Users Authenticating The XEN hypervisor hosting Domain Controller HSU-ADDC03 stopped responding. Authentication should still succeed via two other Active Directory Domain controllers with slight delay. ITS-Sysadmins are investigating. ITS Sysadmins had to reboot other XEN cluster nodes to troubleshoot this problem. XEN hypervisor OS file system ran out of space.
Wireless Network Jan 28 2012 - 10:00 All campus community using HSUWireless and ResnetWireless, HSU Marine Lab, Walter Warren House 38, HSU Corpyard On January 9th, 2012 the local wireless controller stopped reporting information via a non-critical process used in collecting usage statistics. These statistics are important in capacity planning and decision making and being unavailable ITS cannot determine how many IP addresses are available in our wireless pool. The vendor has been unable to resolve the issue in production and is recommending an OS upgrade.
CMS/Peoplesoft Jan 25 2012 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions MP 3.00; MP 3.01; Baseline (requires COMR), College America.
CMS/Peoplesoft Jan 25 2012 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions MP 3.00; MP 3.01; Baseline (requires COMR),College America.
CMS/Peoplesoft Jan 25 2012 - 18:00 Anyone who uses CFS (Finance), CFS Data Warehouse, or its database services CMS Central will install a variety of changes*, none of which affect campus end-users (except for the system outage). *ACH Payment Processing, Investment Distribution Allocation (IDA), Increase Instance Count, and Validate Inter/IntraUnit Setup.
OBI Jan 14 2012 - 09:00 OBI Users The Clone process has finished, OBIPROD is back up.
Other Jan 15 2012 - 05:00 All Users Monthly Maintenance Window
Campus Data Network Jan 19 2012 - 16:00 All Students, Faculty, & Staff occupying the Buck House #97. As part of the campus ITRP2 network project, network equipment will be replaced every 3-5 years to keep the campus data network modern, secure and robust.
Moodle Jan 5 2012 - 20:00 All Moodle users. Space has been added to the Moodle File System.
OBI Jan 5 2012 - 09:36 Finance, HR, Student Oracle database reached 100% capasity. Additional space will be added.
Other Jan 4 2012 - 16:06 person web pages
CMS/Peoplesoft Jan 4 2012 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions MP 2.04, GL Mapping (SFS).
Campus Data Network Jan 12 2012 - 16:00 All Students, Faculty, & Staff occupying the Jensen House #94. As part of the campus ITRP2 network project, network equipment will be replaced every 3-5 years to keep the campus data network modern, secure and robust.
Campus Data Network Jan 5 2012 - 16:00 All Students, Faculty, & Staff occupying the Walter Warren House #38. As part of the campus ITRP2 network project, network equipment will be replaced every 3-5 years to keep the campus data network modern, secure and robust.
Campus Data Network Dec 22 2011 - 23:59 Network users in Art A, Art B, Press Box, Forbes C, Jenkins Hall, Music B, Natural History Museum, Science C, and Van Matre Hall. Emergency Maintenance is needed to reboot a switch that is causing Performance issues resulting from hung processes on Van Matre Hall core switch.
feedback